40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  CTN Lab elog  Not logged in ELOG logo
Message ID: 1508     Entry time: Mon Sep 8 12:41:32 2014
Author: Evan 
Type: DailyProgress 
Category: TempCtrl 
Subject: Chamber temp change 

The beat is currently at 450 kHz. So I've changed the setpoint on the chamber temperature from 31.200 °C to 31.300 °C. We'll see if this pushes the beat to a higher frequency.

Edit: actually I decided to leave the chamber temperature alone and instead adjust the north cavity heater again. With a DVM, I measured the initial heater voltage as 10.64 V. I turned the power supply knob ever so slightly to get 10.60 V instead.

ELOG V3.1.3-